# Improving semantic understanding in speech language models via brain-tuning

### v2

@@ 4.2 Effect of Low-level Features on Brain Alignment
We observe that brain-tuning reduces the impact of low-level features on alignment with late language regions across all model families, including Whisper, which did not originally show improvements in the total brain alignment in these regions (Fig. 2a). The comparison models once again do not account for the improvements due to brain-tuning. Similarly to before, for all models, brain-tuning leads to no significant changes related to the primary auditory regions. These observations also hold on the voxel-level (Fig. 3(c) for Wav2Vec2.0 and one representative participant; see Appendix D.2 for the remaining participants) ). Moreover, brain-tuning can also make up for scale: we observe that a brain-tuned smaller model (HuBERT small) is similarly aligned with late language regions as a substantially larger pretrained model of the same family (HuBERT large), while having lower impact of low-level features on this alignment (see Appendix A.2 for more details).
@@
Figure 4: Downstream task performance for different models. Brain-tuned models’ performance is the mean and STE across participants. Brain-tuned models show consistent improvement over the baselines, with biggest gains in more semantic tasks (ASR and phonetic sentence-type prediction).
@@ 4.3 Downstream Performance
Performance on downstream tasks.
Fig. 4 presents the models’ performances on popular downstream tasks that require different levels of semantic understanding. The brain-tuned performance is averaged across layers and participants, with the exception of ASR where the performance is only averaged across participants since evaluations are done only once per model. We observe that brain-tuning boosts the performance of all model families across all downstream tasks with varying semantic difficulty, with the exception of emotion prediction, in which the performance improves slightly for one model family and decreases slightly for the other two (Fig. 4(f)). Specifically, brain-tuning leads to the biggest gains for tasks that are most semantically challenging and require recognition and understanding beyond the local context. Brain-tuning increases the ASR accuracy on English ASR by up to 12% (Fig. 4(a)), enhances the phonetic sentence type understanding by 20-23% (Fig. 4(b)), and nearly doubles the performance on sequence understanding for the self-supervised model families (Fig. 4(c)). Tasks that relate to local word or phoneme understanding see moderate gains (6% gain in phonemes prediction across model families (Fig. 4(d)), and 3% gain in word identity prediction for the self-supervised models and 17% for Whisper (Fig. 4(e))).
General trends in downstream performance. We observe a few trends in all downstream tasks: (1) the BigSLM fine-tuned models are the same or worse than the pretrained ones, except for ASR, for which BigSLM offers a noticeable improvement over pretraining. (2) Generally, the pretrained, brain-tuned, and BigSLM fine-tuned models perform better than the Random-fMRI baselines, which indicates that permuting fMRI targets does not lead to improvement on downstream tasks. (3) The models perform better than naive classifiers, so substantial gains in performance cannot be attributed to random behavior. (4) More interestingly, the brain-tuned Whisper encoder consistently and substantially outperforms its pretrained and BigSLM-finetuned versions, which often have close to random or naive performances. This shows that the substantial gain in performance is due to fine-tuning with the matched fMRI data.
Change in semantic-phonetic preference. Finally, when we compare the semantic-phonetic preference for the brain-tuned vs. pretrained models (Appendix Fig. 8), we find that the preference becomes more semantic in the late layers for the brain-tuned models. In contrast, the pretrained models show either no change or a decrease in semantic preferences. This indicates that the brain-tuned models have reduced phonetic preference in the later layers. We elaborate on these findings in Appendix C.
Excitingly, the consistent improvement in downstream tasks and reduced phonetic preference in late layers is in tandem with the improvement in brain alignment with late language regions and the reduced impact of low-level features on this brain alignment. These results provide converging evidence for improved semantics in the brain-tuned models.
@@ 5 Discussion
First, two of the three tested model families have improved alignment with new brain recordings in the semantic language regions after brain-tuning (Fig. 2). For all model families, brain-tuning also significantly reduces the impact of low-level speech features on alignment with late language regions (Fig. 3). This is a marked improvement on pretrained models, which were shown to almost entirely rely on low-level speech features to align with semantic brain regions Oota et al. (2024a). Furthermore, brain-tuned models align with semantic language regions as well as larger pretrained models while also having a lower low-level feature impact (Appendix Fig. 5). This suggests that brain-tuning has the potential to break the brain alignment plateau for speech models, which has been shown to reach saturation around 700M parameters (Antonello et al., 2024). Brain-tuning larger models may break this saturation and lead to substantially more alignment in the semantic language regions, further increasing their utility as model organisms (Toneva, 2021).
Second, the brain-tuned models generally improve in performance on a range of downstream tasks with different semantic difficulties (Fig. 4). This shows that brain-tuned models have improved general capabilities and that the fine-tuning process did not degrade the performance or make the models gain evident capabilities in one task but not the others. Moreover, specifically Whisper gains substantially from brain-tuning, improving from close to naive/random F1 scores to significant enhancement across the board. Lastly, the brain-tuned models also improve their semantic preference in the late layers (Appendix Fig. 8). We further emphasize that the gain in performance and semantic preference is due to additional training data of less than $0.7\%$ of the original training size, which is comparable to, if not less than, a typical fine-tuning dataset for a specific task. Yet, in the brain-tuning case, there is an evident gain in performance and a wider generalization across multiple downstream semantic tasks. Overall, to the extent of our knowledge, this is the first work to show substantial downstream performance gains when augmenting an ML model with brain-relevant data, across not only the speech domain but also language and vision.
@@ 6 Conclusion
Reproducibility Statement. In this paper, we fully describe our proposed Brain-tuning approach and how to evaluate it. (1) Sections 3.1, 3.2, and 3.3 detail the exact model families, data processing, and fine-tuning settings and hyper-parameters needed to carry out Brain-tuning for any given model. (2) Section 3.4, Appendix B, and Appendix C detail all evaluation pipelines and metrics, alongside any additional datasets or training settings and hyper-parameters needed for evaluation. (3) We will make the code for brain-tuning and its evaluation publically available once the paper is accepted.
@@ A.2 Brain-tuning Alignment comparison with bigger models
A.3 Brain Alignment With Random-fMRI
We extend Fig. 6 by showing the Random-fMRI baseline alongside the brain-tuned, pretrained, and BigSLM fine-tuned models. Fig. 6(b) clearly shows that Random-fMRI models have the lowest brain alignment values in late language regions compared to even BigSLM fine-tuned baselines, and in the primary auditory regions, it’s also much lower than the pretrained version. This strongly indicates that randomly permuting the fMRI targets while brain-tuning the models with the same stimuli substantially harms the model’s alignment. Hence, having the correct fMRI targets is essential for the alignment results we get from the brain-tuned models.
@@ A.3 Brain Alignment With Random-fMRI
Figure 6: (a), (b) Normalized brain alignment across participants and models for different ROIs, where * means brain-tuned has alignment values that are statistically significantly different from pretrained. Random-fMRI fine-tuned baselines have much lower brain alignment than pretrained models in that late language and primary auditory areas.
A.4 Low-level Impact With Random-fMRI
We extend Fig. 3 by showing the Random-fMRI baseline alongside the brain-tuned, pretrained, and BigSLM fine-tuned models. Fig. 7 clearly shows that Random-fMRI models have the highest low-level impact (highest drop) compared to even BigSLM fine-tuned baselines in both late language regions and the primary auditory region. Even though the original normalized alignment values for the Random-fMRI models are very low to begin with, they still undergo a substantial drop after we remove the low-level features. Similar to the results from Fig. 6, these results also indicate that randomly permuting the fMRI targets while brain-tuning the models with the same stimuli substantially harms the model’s semantics. Hence, having the correct fMRI targets is essential for reduced low-level impact we get in the brain-tuned models.
@@ A.4 Low-level Impact With Random-fMRI
Figure 7: (a), (b) Low-level Speech Features Impact (Percentage Drop in Brain Alignment) for different ROIs, where * means brain-tuned has alignment values that are statistically significantly different from pretrained. Random-fMRI fine-tuned baselines have a much higher low-level impact than pretrained models in that late language and primary auditory areas.
@@ Appendix C Measuring Semantic-Phonetic preference
Finally, when we investigate the semantic-phonetic preference $d$ (the delta of the distances between semantic and phonetic features) in the brain-tuned vs. pretrained models (Fig. 8), we find that $d$ decreases in the late layers for the brain-tuned models, but it either increases or stays the same for the pretrained ones. Our findings about the pretrained model replicate those of Choi et al. (2024). The findings about the brain-tuned models indicate that, unlike the pretrained models, semantic pairs in the late layers of these models are closer to the phonetic ones, and hence the layers are less phonetically dominated (i.e., the semantic preference increases and the phonetic vs semantic preference decreases in these layers). This together with the downstream results from Fig. 4 indicate that the brain-tuned models are not phonetically biased (in late layers) and that they are capable of performing better on both phonetic and semantic tasks.
Appendix D Brain Plots for Different Participants

### v3

@@ 4.2 Effect of Low-level Features on Brain Alignment
We observe that brain-tuning reduces the impact of low-level features on alignment with late language regions across all model families, including Whisper, which did not originally show improvements in the total brain alignment in these regions (Fig.2a). The comparison models once again do not account for the improvements due to brain-tuning. Similarly to before, for all models, brain-tuning leads to no significant changes related to the primary auditory regions. These observations also hold on the voxel-level (Fig.3(c) for Wav2Vec2.0 and one representative participant; see Appendix F.3 for the remaining participants). Moreover, brain-tuning can also make up for scale: we observe that a brain-tuned smaller model (HuBERT small) is similarly aligned with late language regions as a substantially larger pretrained model of the same family (HuBERT large), while having lower impact of low-level features on this alignment (see Appendix A.2 for more details).
@@
Figure 4: Downstream task performance for different models. Brain-tuned models’ performance is the mean and STE across participants. Brain-tuned models show consistent improvement over the baselines, with the biggest gains in more semantic tasks.
@@ 4.3 Downstream Performance
Performance on downstream tasks. Fig.4 presents the models’ performances on popular downstream tasks that require different levels of semantic understanding. The brain-tuned performance is averaged across layers and participants, except for ASR where the performance is only averaged across participants since evaluations are done only once per model. We observe that brain-tuning boosts the performance of the pretrained models for all model families across all tested downstream tasks, with the exception of emotion prediction, in which the performance improves slightly for one model family and decreases slightly for the other two (Fig.4(f)). The brain-tuned models also generally outperform the BigSLM-tuned baselines, with the exception of the emotion prediction task. Most notably, brain-tuning leads to the most consistent gains above the BigSLM-tuned and pretrained baselines for those tasks where fine-tuning with representations from text-based language models is also helpful (Fig.10; Appendix D.2). This provides additional evidence that brain-tuning is most beneficial for tasks where improved semantic understanding is helpful.
General trends in downstream performance. We observe a few trends in all downstream tasks: (1) the BigSLM-tuned models perform on par or better than the pretrained ones, indicating that the BigSLM-tuned model is a meaningful baseline. (2) Generally, the pretrained, brain-tuned, and BigSLM-tuned models perform better than the Random Brain-tuned baselines, which indicates that permuting fMRI targets does not lead to improvement on downstream tasks. (3) The models perform better than naive classifiers, so substantial gains in performance cannot be attributed to random behavior. (4) More interestingly, the brain-tuned Whisper encoder consistently and substantially outperforms its pretrained and BigSLM-tuned versions, which often have close to random or naive performances. This shows that the substantial gain in performance is due to fine-tuning with the matched fMRI data.
Change in semantic-phonetic preference. Finally, when we compare the semantic-phonetic preference for the brain-tuned vs. pretrained models (Fig.8), we find that the preference becomes more semantic in the late layers for the brain-tuned models. In contrast, the pretrained models show either no change or a decrease in semantic preferences. This indicates that the brain-tuned models have reduced phonetic preference in the later layers. We elaborate on these findings in Appendix C.
Excitingly, the consistent improvement in downstream semantic tasks and reduced phonetic preference in late layers is in tandem with the improvement in brain alignment with late language regions and the reduced impact of low-level features on this brain alignment. These results provide converging evidence for improved semantics in the brain-tuned models.
@@ 5 Discussion
First, two of the three tested model families have improved alignment with new brain recordings in the semantic language regions after brain-tuning (Fig.2). For all model families, brain-tuning also significantly reduces the impact of low-level speech features on alignment with late language regions (Fig.3). This is a marked improvement over pretrained models, which were shown to almost entirely rely on low-level speech features to align with semantic brain regions (Oota et al., 2024a). Furthermore, brain-tuned models align with semantic language regions as well as larger pretrained models while also having a lower low-level feature impact (Appendix Fig.5). This suggests that brain-tuning has the potential to break the brain alignment plateau for speech models, which has been shown to reach saturation around 700M parameters (Antonello et al., 2024). Brain-tuning larger models may break this saturation and lead to substantially more alignment in the semantic language regions, further increasing their utility as model organisms (Toneva, 2021).
Second, the brain-tuned models most substantially improve performance on downstream semantic tasks while also maintaining performance on core speech low-level tasks (Fig.4 & Fig.9). This provides additional evidence that brain-tuning improves semantic understanding in speech models while not negatively impacting their fundamental speech capabilities. We attribute the brain-tuned models’ balanced semantic and low-level understanding to the inclusion of voxels from both late language areas and auditory cortex during brain-tuning. We expect that if brain-tuning only considered auditory cortex voxels, there would not be improvements in the model’s semantic understanding. Conversely, if brain-tuning only considered late language areas voxels, the model’s speech performance would substantially degrade. Future work can investigate these hypotheses.
Lastly, the brain-tuned models also improve their semantic preference in the late layers (Appendix Fig.8). We further emphasize that the gain in performance and semantic preference is due to additional training data of less than $0.7\%$ of the original training size, which is comparable to, if not less than, a typical fine-tuning dataset for a specific task. Yet, in the brain-tuning case, there is an evident gain in performance and a wider generalization across multiple downstream semantic tasks. There is also good evidence that expanding brain data and model families further will increase the gains from brain-tuning (Appendix E.2 & E.1). Overall, to the extent of our knowledge, this is the first work to show substantial downstream performance gains when augmenting an ML model with brain-relevant data, across not only the speech domain but also language and vision.
@@ 6 Conclusion
Reproducibility Statement
In this paper, we fully describe our proposed brain-tuning approach and how to evaluate it. (1) Sections 3.1, 3.2, and 3.3 detail the exact model families, data processing, and fine-tuning settings and hyper-parameters needed to carry out Brain-tuning for any given model. (2) Section 3.4, Appendix B, Appendix C, and Appendix D.2 detail all evaluation pipelines and metrics, alongside any additional datasets or training settings and hyper-parameters needed for evaluation. (3) We will make the code for brain-tuning and its evaluation publicly available once the paper is accepted.
Acknowledgments
This work was partially funded by the German Research Foundation (DFG) - DFG Research Unit FOR 5368 and by the CS@max planck graduate center.
@@ A.2 Brain-tuning Alignment comparison with bigger models
A.3 Low-level Speech Features Details
Here, we detail the definition and acquisition method of the low-level speech features used in our computation of low-level impact (namely Power Spectrum, Di-Phones & Tri-Phones, and Articulation ). We obtain these features for the fMRI stories used for brain alignment computation (Section 3.4.1).
Power Spectrum. We follow the method described in (Gong et al., 2023). For each TR time (a 2s segment), we quantify the time-varying power spectrum across 448 frequency bands. This power spectrum is obtained by estimating the power of the sound signal between 25 Hz and 15 kHz, in 33.5 Hz bands (giving a total of 448 bands). For our experiments, we use the power spectrum features from (Deniz et al., 2019).
Di-Phones & Tri-Phones. For a given speech utterance, DiPhones represent the adjacent pair of phones (e.g., [da], [aI]). For each audio segment of length 2 seconds (TR length), we have a one-hot encoding vector representing the presence or absence of all possible 858 diphones. The same is done for Tri-Phones but for triplets of phones instead of pairs. The annotations of this data were done using the Praat software and were obtained from the shared data by the authors of (Oota et al., 2024a). [^1] https://www.fon.hum.uva.nl/praat/
Articulation. Phoneme articulations are the specific ways speech organs (e.g., tongue or vocal cord) move to produce different phonemes. They associate each phoneme with a set of properties that specify things like whether this phoneme makes vocal cords vibrate. We use phoneme articulations as mid-level speech features, mapping hand-labeled phonemes to a set of 22 articulatory characteristics. For our experiments, we use the articulation annotations from (Deniz et al., 2019).
A.4 Brain Alignment With Random Brain-tuned
We extend Fig.6 by showing the Random Brain-tuned baseline alongside the brain-tuned, pretrained, and BigSLM-tuned models. Fig.6(b) clearly shows that Random Brain-tuned models have the lowest brain alignment values in late language regions compared to even BigSLM-tuned baselines, and in the primary auditory regions, it’s also much lower than the pretrained version. This strongly indicates that randomly permuting the fMRI targets while brain-tuning the models with the same stimuli substantially harms the model’s alignment. Hence, having the correct fMRI targets is essential for the alignment results we get from the brain-tuned models.
@@ A.4 Brain Alignment With Random Brain-tuned
Figure 6: (a), (b) Normalized brain alignment across participants and models for different ROIs, where * means brain-tuned has alignment values that are statistically significantly different from pretrained. Random Brain-tuned baselines have much lower brain alignment than pretrained models in the late language and primary auditory areas.
A.5 Low-level Impact With Random Brain-tuned
We extend Fig.3 by showing the Random Brain-tuned baseline alongside the brain-tuned, pretrained, and BigSLM-tuned models. Fig.7 clearly shows that Random Brain-tuned models have the highest low-level impact (highest drop) compared to even BigSLM-tuned baselines in both late language regions and the primary auditory region. Even though the original normalized alignment values for the Random Brain-tuned models are very low to begin with, they still undergo a substantial drop after we remove the low-level features. Similar to the results from Fig.6, these results also indicate that randomly permuting the fMRI targets while brain-tuning the models with the same stimuli substantially harms the model’s semantics. Hence, having the correct fMRI targets is essential for reduced low-level impact we get in the brain-tuned models.
@@ A.5 Low-level Impact With Random Brain-tuned
Figure 7: (a), (b) Low-level Speech Features Impact (Percentage Drop in Brain Alignment) for different ROIs, where * means brain-tuned has alignment values that are statistically significantly different from pretrained. RandomBrain-tuned baselines have a much higher low-level impact than pretrained models in that late language and primary auditory areas.
@@ Appendix C Measuring Semantic-Phonetic preference
Finally, when we investigate the semantic-phonetic preference $d$ (the delta of the distances between semantic and phonetic features) in the brain-tuned vs. pretrained models (Fig.8), we find that $d$ decreases in the late layers for the brain-tuned models, but it either increases or stays the same for the pretrained ones. Our findings about the pretrained model replicate those of (Choi et al., 2024). The findings about the brain-tuned models indicate that, unlike the pretrained models, semantic pairs in the late layers of these models are closer to the phonetic ones, and hence the layers are less phonetically dominated (i.e., the semantic preference increases and the phonetic vs semantic preference decreases in these layers). This, together with the downstream results from Fig.4 indicates that the brain-tuned models are not phonetically biased (in late layers) and that they are capable of performing better on both phonetic and semantic tasks.
Appendix D Additional Tasks and Baselines
In this section, we extend the evaluation of brain-tuned models to include two low-level tasks, namely MFCC and FBANK prediction. We also add language model baseline (comparison model) and report its downstream and low-level performance relative to to the other models (i.e., brain-tuned, Pretrained and BigSLM-tuned).
D.1 Brain-tuned low-level performance
For MFCC prediction, we train linear probes to predict the MFCC coefficients, while for FBANK prediction, we train linear probes to predict the filter banks. For both tasks, we use the TIMIT dataset (Garofolo, 1993) and we evaluate using the $R^{2}$ coefficient, averaged across layers.
Fig.9 reports pretrained, BigSLM-tuned, and brain-tuned models’ performance on two low-level tasks (MFCC prediction and FBANK prediction); it shows that brain-tuned and BigSLM-tuned models don’t perform substantially lower than their pretrained counterparts. This indicates that fine-tuned models are still capable of doing these core low-level tasks just as well as their pretrained version. We think this evidence reduces the possibility that catastrophic forgetting happened after fine-tuning.
(a) MFCC prediction mean $R^{2}$ values
(b) FBANK prediction mean $R^{2}$ values
(c) $R^{2}$ values of MFCC prediction layer-wise
(d) $R^{2}$ values of FBANK prediction layer-wise
Figure 9: Low-level task performance comparison for different models. It signifies that brain-tuned models didn’t lose low-level capabilities at the expense of gaining more semantics, as the averaged $R^{2}$ values are very close and the layer-wise values are the same in early layers and slightly decreased in late layers.
D.2 Details for Comparison Baselines
We elaborate here the comparison models introduced in Section 3.3.1 and provide a complete contrast for all comparison models on all downstream tasks for the Wav2Vec2.0 model family.
Random brain-tuned. In order to test how the addition of any fMRI data impacts model performance, this baseline uses the same fine-tuning process for brain-tuning (Section 3.3) but with incorrect randomly chosen fMRI targets. Instead of using the correct fMRI responses for the input stimulus, we use block-permuted fMRI responses that correspond to different stimuli. We carry out the fine-tuning process for each participant, and the training setup is identical to the brain-tuning ones (i.e., the same learning rates, learning rate schedule, optimizer, and number of epochs). The same set of voxels used in brain-tuning is also used for this baseline.
Big spoken language model-tuned (BigSLM-tuned). We aim to test the importance of having fMRI responses as the fine-tuning targets with this baseline. To this end, we replace the fMRI target vectors for the input stimuli with representations for the same stimuli obtained from the BigSLM layers. This creates a proxy for the fMRI targets that is rich in meaningful spoken-language related information. For this baseline, we utilize Whisper Medium (800M parameters) as the BigSLM and use a concatenation of all its decoder layers’ representations. The fine-tuning objective is then to construct the concatenated representations vectors from the BigSLM layers (instead of reconstructing the fMRI responses). We use the same training setup and the same stimulus data that were used for brain-tuning (Section 3.3).
Stimulus-tuned. This baseline tests whether tuning with the fMRI signal results in additional gains over simply further tuning using the stimulus audio only. This highlights any improvements in the model performance that would be only due to training with more data. To isolate the effect of the audio stimulus alone, we develop a model whose pre-training includes the audio stimulus used in the fMRI data (LeBel et al., 2024). To this end, for additional pre-training on the stimulus audio data, we use the same losses with the same weights of Wav2Vec2.0 self-supervised pre-training (the diversity loss and the contrastive loss) (Baevski et al., 2020). This stimulus-tuned model is trained for 200 epochs and uses a base learning rate of $2\times 10^{-5}$ with a warm-up for the first 10% of the updates and then a linear decay schedule.
Text language model-tuned (LM-tuned). We develop two text language model-tuned baselines to test the potential of using a text language model (LM) in a similar fashion to the BigSLM-tuned baseline. The first one uses representations from the pretrained GPT2-Medium model (Radford et al., 2019) which is similar in size to the BigSLM model, and the second one uses representations from Llama2-7B Parameter Model (Touvron et al., 2023) which is a much bigger and more recent large LM. We apply a similar fine-tuning pipeline of brain-tuning with the concatenated representations of the corresponding LM as targets. For the GPT2-Medium model, we concatenate the representations of all layers, while for LLama2-7B one in every 3 layers is taken. We found that LM-tuned models need more training than the BigSLM-tuned one; we train them for 200 epochs. We report the downstream performance of these baselines (the GPT2-tuned and the Llama2-tuned) for the Wav2Vec2.0 model family.
The stimulus-tuned model performs very similarly to the pretrained model, indicating that further self-supervised training on these additional few hours of stimulus data doesn’t impact downstream model performance (Fig.10). As for the LM-tuned baselines, GPT2-tuned and Llama2-tuned models improve over the pretrained and BigSLM-tuned ones on two tasks (namely Phonemes and Phonetic Sentence Type Prediction), with the Llama2-tuned being slightly better than the brain-tuned one on the same two tasks. The jump from pretrained to GPT2-tuned to Llama2-tuned indicates that more semantics lead to better performance on Phonemes and Phonetic Sentence Type Prediction tasks. These tasks are also the ones for which brain-tuning leads to the largest improvement over the pretrained models, further supporting our conclusions that brain-tuning can improve semantic understanding in speech models. The results for the Random Brain-tuned and BigSLM-tuned baselines are identical to Fig.4; they mainly emphasize that Random Brain-tuning hurts the model performance and BigSLM-tuning doesn’t help strongly with the more semantic tasks.
Figure 10: Downstream task performance of brain-tuned models vs. all baselines for Wav2Vec2.0. The additional baselines beyond those presented in Fig.4 are shown in blue hues. We observe that LM-tuned baselines (GPT2-tuned and Llama2-tuned) improve over the pretrained model on tasks where we see big improvements with brain-tuning, further supporting our conclusions that brain-tuning can improve semantic understanding in speech models. The stimulus-tuned model performs very similarly to the pretrained one, indicating that simply further training on the audio stimulus is not sufficient for the gains observed from brain-tuning.
Appendix E More studies on downstream tasks
E.1 Downstream performance comparison with HuBERT Large
To complement our results on the brain alignment gap between HuBERT Large and the Brain-tuned Base version (Appendix A.2), we report here the gap between them in downstream performance (reporting the same participants for brain-tuned models). Fig.11 shows that for HuBERT, the brain-tuned base architecture is closing on the pretrained large architecture. In other words, we gain a performance close to the large architecture with the same number of parameters of the base architecture (after it’s brain-tuned).
Figure 11: Downstream performance comparison of pretrained HuBERT large architecture vs Brain-tuned HuBERT base architecture. The HuBERT base architecture performs closely to the large pretrained architecture, bridging the gap between the base and large model sizes.
(a) Performance Improvement vs. Model Size for HuBERT
(b) Performance Improvement vs. Data Size for Wav2Vec2.0
Figure 12: Performance Improvement (brain-tuned -- pretrained performance) across model sizes and data sizes. We see clearly that lowering the amount of data limits the gain in performance due to brain-tuning, and the same goes for increasing the model size while keeping the data fixed.
E.2 Effect of Model and data sizes on Brain-tuning
We study here how brain-tuning is affected when we increase the pretrained model size or decrease the size of the data used for brain-tuning. We do this by plotting the performance improvement with brain-tuning (brain-tuning performance -- pretrained performance) across model sizes or data sizes, for a range of downstream tasks. Fig.12 shows that as the model size increases, the gain in performance due to brain-tuning on downstream tasks drops. When we lower the amount of data used for brain-tuning, the gain in performance due to brain-tuning also drops. This provides a clear signal that data size plays a key role in the positive results we see from brain-tuning.
Appendix F Additional Brain Plots for Different Participants